OpenStudy (gokuporter):

Checking my answer. Write the equation on the line in Slope-Intercept form. Point on the line: (-5,-4) & (3,-7) Slope: -3/8 Slope Intercept form: y+4= -3/8(x+5) Answer: y= 3/8x + 17/8

OpenStudy (gokuporter):

@dan815 Is this correct

OpenStudy (gokuporter):

@pooja195

hartnn (hartnn):

y+4= -3/8(x+5) The next step would be y = -3x/8 - 15/8 -4 right??

hartnn (hartnn):

the only error is: there should be -3x/8 instead of just 3x/8

OpenStudy (gokuporter):

Typo

hartnn (hartnn):

ohk, then y= -3/8x + 17/8 is correct! good :)

OpenStudy (gokuporter):

:)